OpenStudy (anonymous):

Write the equation in standard form for the line that passes through the points ( - 5, 1 ) and ( 6, -2 ).

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -5}}\quad ,&{\color{blue}{ 1}})\quad &({\color{red}{ 6}}\quad ,&{\color{blue}{ -2}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and move everything to the left}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

y=mx+b do y2-y1/x2-xiget slope and use a y value for b mx is slope

OpenStudy (anonymous):

Standard Form Is Ax+By=C

OpenStudy (anonymous):

3x+11y=-4 is the answer

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np bye